OpenStudy (anonymous):

write the equation of the line that is parallel to the line 2x+3y=5 and contains the point (-2 1)

OpenStudy (jdoe0001):

hmmm if you solve "2x+3y=5" for "y"..... what would it look like?

OpenStudy (texaschic101):

basically what jdoe is saying, is in y = mx + b form, the slope is in the m position. So solve for y is the same thing

OpenStudy (anonymous):

y=-2x/3+5/3?

OpenStudy (texaschic101):

2x + 3y = 5 3y = -2x + 5 y = -2/3x + 5/3 y = mx + b ....slope is in the m position so what is the slope ?

OpenStudy (anonymous):

-2/3

OpenStudy (jdoe0001):

so the slope of a line parallel to that one, will have the same slope as that one... thus \(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ 1}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= -\cfrac{2}{3} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (texaschic101):

correct....and a parallel line will have the same slope. so now use y = mx + b again slope(m) = -2/3 (-2,1).....x = -2 and y = 1 we already know the slope(m), but we need to find the y intercept(b) so we sub y = mx + b 1 = -2/3(-2) + b 1 = 4/3 + b 1 - 4/3 = b 3/3 - 4/3 = b -1/3 = b so your equation in slope intercept form is : y = -2/3x - 1/3 do you have any questions ?

OpenStudy (jdoe0001):

well..... you're not asked to put it in slope-intercept form, so.... I gather you don't have to solve it for "y"

OpenStudy (anonymous):

no questions

OpenStudy (texaschic101):

good to hear :)....just remember, that a parallel line will have the same slopes, but perpendicular lines are another story

OpenStudy (anonymous):

thanks to both :P